# Are Dense Labels Always Necessary for 3D Object Detection from Point Cloud?

### v1

@@ 5.3 Comparisons with Fully-Supervised Methods
[…]
@@
500* scenes + ${}^{\dagger}$ instances WS3D [24] 85.04 75.94 74.38 88.95 85.83 85.03 82.52
${}^{\dagger}$ instances Ours (PointRCNN-based [14]) 84.69 71.93 67.06 89.57 84.07 78.81 79.36
[…]
@@
5.4 Comparisons with Weakly-Supervised Methods
Comparison with the semi-supervised method. We compare the proposed SS3D+ method with semi-supervised method 3DIoUMatch [11], which is based on the advanced detector PV-RCNN [6]. To make a fair comparison, we also adopt the PV-RCNN as the base detector and keep all methods trained with the same number of annotated objects. Specifically, in the KITTI train split, there are 3,712 scenes and these scenes contain a total of 17,289 objects for cars, pedestrians, and cyclists. For semi-supervised methods, 1% labeled data means 37 ($3712\times 1\%$) scenes, which include an average of 172 ($17289\times 1\%$) labeled objects used for training. So as for 1% labeled data in our extremely sparse split, we randomly select 172 scenes including 172 labeled objects for training. We also test the case of 2% labeled training data for both methods. Generally, our sparse annotation strategy may provide an easier way to annotate objects from point cloud scenes compared to a dense annotation strategy for semi-supervised methods. The detection results with different ratios of labeled data can be seen in Fig. 11. As we can see, our SS3D++ (the green bar) significantly outperforms the current SOTA semi-supervised method, 3DIoUMatch, for three different classes with all ratios of labeled training data split. Fig. 12 illustrates the comparison of mining pseudo instances by 3DIoUMatch and SS3D++ methods. The 3DIoUMatch method may easily fall into a suboptimal solution due to limited information transfer from labeled to large discrepancy unlabeled scenes, while the SS3D++ method can encourage easier intra-scene information transfer due to the proposed sparse annotation strategy, in which each scene has one object annotation. In fact, the results of the green bar in Fig. 11 only use 172 (or 344) scenes for training, while 3DIoUMatch has access to all 3712 scenes involved in training. This illustrates the potential of the SS3D++, and we can further boost the performance of the SS3D++ by utilizing the remaining unlabeled scenes, and obtain more improvements as depicted by the yellow bar in Fig. 11, outperforming 3DIoUMatch method by a large margin (more details can refer to Section 5.6).
[…]
@@ 5.5 Ablation Study and Analysis
In this section, we present a series of ablation studies to analyze the effects of different modules in our proposed SS3D++ method. Following the aforementioned setting, all models are trained on the KITTI dataset with an extremely sparse split and evaluated on the val split. We take PointRCNN[14] as an example base detector to conduct our ablation study, and our methods can obtain similar results for other detectors. Table V summarizes the ablation results on our reliable background mining module (RBMM) and confident missing-annotated instances mining module (CMIM). All results are reported by the mAP with 40 recall positions under moderate difficulty.
Effect of the reliable background mining module. In the $1^{st}$ row of Table V, we discarded both modules, so it represents the standard PointRCNN detector trained with the extremely sparse split, whose performances have significant degradation compared with full annotations. In the $2^{nd}$ row, we leverage the reliable background mining module to extract reliable backgrounds. It can be observed that the performance of the three categories is significantly improved, especially for the ”car” category, which has shown a substantial increase. This is due to that the autonomous driving dataset usually contains the largest number of cars. Therefore, under the sparsely-supervised setting, the detector trained with massive missing-annotated instances like cars may severely degenerate detection performance. Fortunately, the proposed reliable background mining module could reduce the risk of treating the missing-annotated instances as background, and thereby facilitate the robust training of the detector by eliminating the negative influence of noise background in the point cloud scenes.
@@
TABLE V: Ablation study of different components for our proposed SS3D++ method. We report the mAP with 40 recall positions. RBMM: reliable background mining module. CMIM: confident missing-annotated instances mining module.
@@
TABLE VI: Ablation study of different components for CMIM of our proposed SS3D++ method. We report the mAP with 40 recall positions. DLF: dynamic loss-based filtering. DCS: dynamic consistency-guided suppression. DCF: density-aware curriculum filtering.
DLF DCS DCF 3D Detection (Mod.) Avg.
@@
Effect of the confident missing-annotated instance mining module. As shown in the $3^{rd}$ row of Table V, the confident missing-annotated instance mining module can also improve the baseline by mining high-quality instances. We can observe that the performance of the three categories is significantly improved, especially for the “pedestrian” and “cyclist” categories, which have shown better improvement compared with only RBMM. Actually, both of “pedestrian” and “cyclist” categories are relatively hard to detect compared with the ”car” category, naturally leading to more noisy pseudo-annotated instances. Our multi-criteria sample selection process can effectively filter out noisy pseudo-annotated instances, and thereby brings more robust detection performance. Therefore, it could bring mutual amelioration when we combine CMIM and RBMM ($4^{th}$ row), and achieves better performances than using either of the two.
[…]
@@ 5.6 Learning from Unlabeled Training Scenes
Learning from unlabeled training scenes in an offline manner. In Section 5.4, our SS3D++ method only employs the sparse annotated scenes. Here, we explore to additionally exploit the remaining unlabeled scenes to train the detector. Observing from the yellow bar in Fig. 11 and Fig. 15 (b), we can see that our SS3D++ method can achieve a large improvement by utilizing additional unlabeled scenes, making it superior over weakly-supervised 3DIoUMatch [11] and WS3D [24] methods. Specifically, comparing our method with the 3DIoUMatch [11] in terms of the coverage of generated pseudo instances for unlabeled scenes, we find that 3DIoUMatch achieves only 28% coverage, while our method achieves about 60% coverage. The corresponding accuracy of mined instances in terms of the car, pedestrian, and cyclist can reach 92%, 85%, and 88% for the SS3D++ method. This highlights the potential practical value of exploiting unlabeled scenes. Additional analysis of generating pseudo-annotated instances can be found in Appendix Fig. 3 and Fig. 4. Compared with WS3D [24], our sparse supervision is easy to annotate and universal, making our method capable of generating confident pseudo annotations for easily obtained unlabeled point cloud scenes.
@@
TABLE VII: Annotation analysis of three different sparsely-annotated forms.
Method Data Car - 3D Detection Car - BEV Detection
Easy Mod Hard Easy Mod Hard
Voxel-RCNN[23] Full 92.38 85.29 82.86 95.52 91.25 88.99
Voxel-RCNN[23] (1) Sparse 72.97 63.05 60.23 81.64 74.91 71.93
Ours (1) Sparse 93.50 84.36 82.64 96.63 91.31 89.02
Voxel-RCNN[23] (2) Sparse 88.86 63.04 47.90 93.61 70.82 52.77
Ours (2) Sparse 92.89 83.46 79.81 96.40 91.63 86.73
Voxel-RCNN[23] (3) Sparse 29.15 42.92 49.62 35.81 53.83 60.90
Ours (3) Sparse 94.24 84.11 82.44 96.65 91.38 88.99
5.7 Annotation Analysis
In our sparsely-annotated framework, we just annotate one 3D object per scene. One natural question is to choose which object to annotate. In practice, most annotators may be more inclined to choose the easiest object to annotate, e.g., no occluded objects or high-density objects. To further evaluate the robustness of our SS3D++ method against the hardness of annotated objects, we design the following three different annotated strategies: (1) randomly choose one object from the scene to annotate as default; (2) prioritize choosing one ‘easy’ (close and no truncated) object from the scene; (3) prioritize choosing one ‘hard’ (remote and truncated) object from the scene. Besides, we repeat the experiment 3 times and average the results to eliminate the influence of experimental randomness. From Table VII, we can observe that the hardness of annotated objects has a significant impact on the performance of the fully supervised detector, in which annotated strategy (3) severely deteriorates the detectors. However, by employing the proposed SS3D++ method, regardless of different annotated strategies, such annotation bias can be almost eliminated. Especially, the SS3D++ can achieve on-par performance compared to detectors trained with fully-supervised data for all cases. This shows our method could perform robustly with different annotated strategies, which highlights the capability of our method to effectively address such low-quality annotations. This further substantiates the potential practical value.
@@ 6 Conclusion
Supplementary Material for
”Are Dense Labels Always Necessary for
3D Object Detection from Point Cloud?”
Chenqiang Gao, Chuandong Liu, Jun Shu, Fangcen Liu, Jiang Liu, Luyu Yang, Xinbo Gao, and Deyu Meng
C. Gao is with the School of Intelligent Systems Engineering, the Shenzhen Campus of Sun Yatsen University, Sun Yat-sen University, Shenzhen, Guangdong 518107, China. E-mail: gaochq6@sysu.edu.cn. C. Liu is with the School of Computer Science, Wuhan University, Wuhan 430072, China. E-mail: chuandong.liu@whu.edu.cn. J. Shu and D. Meng are with School of Mathematics and Statistics and Ministry of Education Key Lab of Intelligent Networks and Network Security, Xi’an Jiaotong University, Shaanxi 710049, China. E-mail: xjtushujun@gmail.com, dymeng@mail.xjtu.edu.cn. F. Liu and X. Gao are with School of Communication and Information Engineering, Chongqing University of Posts and Telecommunications, Chongqing 400065, China. E-mail: gaoxb@cqupt.edu.cn, liufc67@gmail.com. J. Liu is with Meta, Menlo Park 94025, USA. E-mail: jiangliu@meta.com. L. Yang is with University of Maryland, College Park 20742, USA. E-mail: loyo@umiacs.umd.edu. Manuscript received April 19, 2005; revised August 26, 2015.
Appendix A Additional Analysis of Visualization Comparison
A.1 Comparison on the KITTI Dataset

### v2

@@ 5.4 Comparisons with SOTA Methods
TABLE VII: Performance comparison on KITTI val split with center-click weakly-supervised WS3D [29] method. We report the mAP with 11 recall positions. ‘*’ denotes the scenes with center-click annotations and $\dagger$}’ denotes precisely-annotated instances.
@@
500* scenes + 534 instances WS3D [29] 85.04 75.94 74.38 88.95 85.83 85.03 82.52
534 instances Ours (PointRCNN-based [44]) 84.69 71.93 67.06 89.57 84.07 78.81 79.36
Ours (PartA2-based [45]) 88.56 77.91 72.46 90.00 86.45 82.28 82.94
@@
Then, We compare the proposed SS3D+ method with semi-supervised method 3DIoUMatch [63] on the KITTI dataset [10]. To make a fair comparison, we also adopt the PV-RCNN as the base detector and keep all methods trained with the same number of annotated objects. Specifically, in the KITTI train split, there are 3,712 scenes and these scenes contain a total of 17,289 objects for cars, pedestrians, and cyclists. For semi-supervised methods, 1% labeled data means 37 ($3712\times 1\%$) scenes, which include an average of 172 ($17289\times 1\%$) labeled objects used for training. So as for 1% labeled data in our extremely sparse split, we randomly select 172 scenes including 172 labeled objects for training. We also test the case of 2% labeled training data for both methods. Generally, our sparse annotation strategy may provide an easier way to annotate objects from point cloud scenes compared to a dense annotation strategy for semi-supervised methods. The detection results with different ratios of labeled data can be seen in Fig. 11. As we can see, our SS3D++ (the green bar) significantly outperforms the current SOTA semi-supervised method, 3DIoUMatch, for three different classes with all ratios of labeled training data split. Fig. 12 illustrates the comparison of mining pseudo instances by 3DIoUMatch and SS3D++ methods. The 3DIoUMatch method may easily fall into a suboptimal solution due to limited information transfer from labeled to large discrepancy unlabeled scenes, while the SS3D++ method can encourage easier intra-scene information transfer due to the proposed sparse annotation strategy, in which each scene has one object annotation. In fact, the results of the green bar in Fig. 11 only use 172 (or 344) scenes for training, while 3DIoUMatch has access to all 3712 scenes involved in training. This illustrates the potential of the SS3D++, and we can further boost the performance of the SS3D++ by utilizing the remaining unlabeled scenes, and obtain more improvements as depicted by the yellow bar in Fig. 11, outperforming 3DIoUMatch method by a large margin (more details can refer to Sec. 5.6).
@@
Comparison with the center-click weakly-supervised method. For the weakly-supervised method, WS3D [29], trains the detector with the first 500 scenes with center-click labels, mixed with randomly selected 534 precisely-annotated instances. Since the standard detectors are not applicable with center-click labels, we only use the same number of (534) precisely-annotated instances randomly chosen from the first 534 scenes to train our proposed SS3D++ method. From Tab. VII, we can easily observe that the SS3D++ method is able to achieve better performance than WS3D for three different base detectors. WS3D may suffer from the significant disparity between point-level weak annotations and 3D precise annotations, which we only have access to precisely-annotated instances, without point-level annotation efforts, which prevents detectors from overfitting the disparity. Moreover, it is worth noting that we can obtain better performance gains when we use more powerful detectors, e.g., from PointRCNN to PartA2. This highlights the advantage of our detector-agnostic framework, hinting at the potential of our method which employs advanced SOTA detectors to achieve competitive performance. Besides, the SS3D++ method can employ additional unlabeled scenes to further boost the performance (refer to Sec. 5.6), while WS3D cannot. This makes our SS3D++ easily apply to real-world 3D detection.
@@ 5.5 Ablation Study and Analysis
In this section, we present a series of ablation studies to analyze the effects of different modules in our proposed SS3D++ method. Following the aforementioned setting, all models are trained on the KITTI dataset with an extremely sparse split and evaluated on the val split. As for the ablation studies on the Waymo dataset, please refer to the Appendix A. We take PointRCNN[44] as an example base detector to conduct our ablation study, and our methods can obtain similar results for other detectors. Tab. VIII summarizes the ablation results on our reliable background mining module (RBMM) and confident missing-annotated instances mining module (CMIM). All results are reported by the mAP with 40 recall positions under moderate difficulty.
Effect of the reliable background mining module. In the $1^{st}$ row of Tab. VIII, we discarded both modules, so it represents the standard PointRCNN detector trained with the extremely sparse split, whose performances have significant degradation compared with full annotations. In the $2^{nd}$ row, we leverage the reliable background mining module to extract reliable backgrounds. It can be observed that the performance of the three categories is significantly improved, especially for the ”car” category, which has shown a substantial increase. This is due to that the autonomous driving dataset usually contains the largest number of cars. Therefore, under the sparsely-supervised setting, the detector trained with massive missing-annotated instances like cars may severely degenerate detection performance. Fortunately, the proposed reliable background mining module could reduce the risk of treating the missing-annotated instances as background, and thereby facilitate the robust training of the detector by eliminating the negative influence of noise background in the point cloud scenes.
@@
TABLE VIII: Ablation study of different components for our proposed SS3D++ method. We report the mAP with 40 recall positions. RBMM: reliable background mining module. CMIM: confident missing-annotated instances mining module.
@@
TABLE IX: Ablation study of different components for CMIM of our proposed SS3D++ method. We report the mAP with 40 recall positions. DLF: dynamic loss-based filtering. DCS: dynamic consistency-guided suppression. DCF: density-aware curriculum filtering.
row DLF DCS DCF 3D Detection (Mod.) Avg.
@@
Effect of the confident missing-annotated instance mining module. As shown in the $3^{rd}$ row of Tab. VIII, the confident missing-annotated instance mining module can also improve the baseline by mining high-quality instances. We can observe that the performance of the three categories is significantly improved, especially for the “pedestrian” and “cyclist” categories, which have shown better improvement compared with only RBMM. Actually, both of “pedestrian” and “cyclist” categories are relatively hard to detect compared with the ”car” category, naturally leading to more noisy pseudo-annotated instances. Our multi-criteria sample selection process can effectively filter out noisy pseudo-annotated instances, and thereby brings more robust detection performance. Therefore, it could bring mutual amelioration when we combine CMIM and RBMM ($4^{th}$ row), and achieves better performances than using either of the two.
To comprehend the roles of different components in CMIM, we further conduct a series of ablation experiments, as shown in Tab. IX. From the 2nd and 3rd row of Tab. IX, we can see that DLF or DCS severely deteriorates the performance of the “pedestrian” category. This is because “pedestrian” category tends to be smallest object in the training scenes, implying that the pseudo annotations are extremely noisy (even under the full supervision, the mAP score is about 60%). Both DLF and DCS depend on the quality of pseudo annotations, and thus their performance may degenerate for the “pedestrian” category with noisy pseudo annotations. While we combine DLF and DCS strategy (the 5th row of Tab. IX), the mAP of the “pedestrian” category reaches +2.23% point compared with the baseline (1st row). As shown in Fig. 13, using DLF or DCS alone leads to a significant number of erroneous pseudo instances for ”pedestrian”. However, combining both components can greatly mitigate this issue.
It can be observed that using the DCF module alone results in negative effects, as it is challenging to rely solely on density as a criterion for judging incorrect pseudo-labels. However, combining the DCF module with either the DLF or DCS modules leads to performance gains. A large performance gain for “pedestrian” (+6% point) can be obtained when we additionally employ the DCF strategy (from the 5th row to the 8th row of Tab. IX). Meanwhile, Fig. 15 shows the quality improvement of pseudo instances by embedding the DCF with the easy-to-hard curriculum manner.
@@ 5.6 Learning from Unlabeled Training Scenes
Learning from unlabeled training scenes in an offline manner. In Sec. 5.4, our SS3D++ method only employs the sparse annotated scenes. Here, we explore to additionally exploit the remaining unlabeled scenes to train the detector. Observing from the yellow bar in Fig. 11 and Fig. 14 (b), we can see that our SS3D++ method can achieve a large improvement by utilizing additional unlabeled scenes, making it superior over weakly-supervised 3DIoUMatch [63] and WS3D [29] methods. Specifically, comparing our method with the 3DIoUMatch [63] in terms of the coverage of generated pseudo instances for unlabeled scenes, we find that 3DIoUMatch achieves only 28% coverage, while our method achieves about 60% coverage. The corresponding accuracy of mined instances in terms of the car, pedestrian, and cyclist can reach 92%, 85%, and 88% for the SS3D++ method. This highlights the potential practical value of exploiting unlabeled scenes. Additional analysis of generating pseudo-annotated instances can be found in Appendix Fig. 20 and Fig. 21. Compared with WS3D [29], our sparse supervision is easy to annotate and universal, making our method capable of generating confident pseudo annotations for easily obtained unlabeled point cloud scenes.
@@ 6 Conclusion
Appendix A Additional experiments on the Waymo dataset
We would like to mention that our method is a general framework, producing consistent ablation experiment results across different datasets, such as KITTI [10] and Waymo [57]. We conduct ablation study on the large-scale Waymo dataset, with Voxel-RCNN [5] as our backbone detector, which shows similar results as on KITTI dataset. Tab. X summarizes the ablation results on our reliable background mining module (RBMM) and confident missing-annotated instances mining module (CMIM). We report the performance of 3D bounding box mean average precision (mAP) and mAP weighted by heading accuracy (mAPH) for objects of two difficulty levels (LEVEL 1 and LEVEL 2). In the $1^{st}$ row of Tab. X, we discarded both modules, so it represents the standard fully-supervised detector trained with the extremely sparse split, whose performances have significant degradation compared with full annotations. The proposed reliable background mining module could reduce the risk of treating the missing-annotated instances as background, and thereby facilitate the training under the sparsely-supervised setting. The confident missing-annotated instance mining module can also improve the baseline by mining high-quality instances, as illustrated in the $3^{rd}$ row of Tab. X. Using the reliable background mining module alone can lead to some improvements in the pedestrian and cyclist categories, but it results in a decline in vehicle performance. This is because, in the Waymo dataset, vehicles are the most numerous objects within scenes and include categories with significant size variations, such as “bus” and “truck”. This variation makes it challenging to achieve high-quality instances mining with ambiguous background distraction.
TABLE X: Ablation study of different components on the Waymo dataset for our proposed SS3D++ method. RBMM: reliable background mining module. CMIM: confident missing-annotated instances mining module.
Veh. (LEVEL 1) Veh. (LEVEL 2) Ped. (LEVEL 1) Ped. (LEVEL 2) Cyc. (LEVEL 1) Cyc. (LEVEL 2)
RBMM CMIM mAP mAPH mAP mAPH mAP mAPH mAP mAPH mAP mAPH mAP mAPH Avg.
- - 47.83 46.32 41.83 40.51 25.88 20.73 21.78 17.44 50.29 48.48 48.38 46.65 38.01
- 63.51 62.92 55.21 54.69 51.32 45.77 43.59 38.83 50.40 49.38 48.47 47.48 50.96
- 42.77 42.38 36.81 36.47 33.94 30.97 28.31 25.82 58.73 57.74 56.47 55.52 42.16
65.99 65.29 57.50 56.90 53.44 48.89 46.41 42.37 60.13 58.93 57.40 56.26 55.79
Tab. XI presents the ablation study results for different components of CMIM within the proposed SS3D++ method, specifically evaluating the impact of Dynamic Loss-based Filtering (DLF), Dynamic Consistency-guided Suppression (DCS), and Density-aware Curriculum Filtering (DCF). When none of the components (DLF, DCS, DCF) are enabled, it represents the model with reliable background mining module. When using DLF, DCS, or DCF alone for pseudo instance mining, the average performance declines to varying degrees. This is because the Waymo dataset contains complex and diverse samples, making it difficult to accurately determine pseudo-labels using a single criterion for different classes. Our multi-criteria sample selection process offers varying degrees of improvement, and when all three criteria are applied together, the greatest benefit is achieved.
TABLE XI: Ablation study of different components for CMIM of our proposed SS3D++ method. DLF: dynamic loss-based filtering. DCS: dynamic consistency-guided suppression. DCF: density-aware curriculum filtering.
Veh. (LEVEL 1) Veh. (LEVEL 2) Ped. (LEVEL 1) Ped. (LEVEL 2) Cyc. (LEVEL 1) Cyc. (LEVEL 2)
DLF DCS DCF mAP mAPH mAP mAPH mAP mAPH mAP mAPH mAP mAPH mAP mAPH Avg.
- - - 63.51 62.92 55.21 54.69 51.32 45.77 43.59 38.83 50.40 49.38 48.47 47.48 50.96
- - 65.70 65.02 57.55 56.94 43.94 39.60 37.58 33.82 50.79 49.84 48.84 47.93 49.79
- - 66.73 65.97 58.55 57.88 38.46 34.76 32.86 29.65 54.00 52.92 51.93 50.89 49.55
- - 60.24 59.17 51.77 50.39 31.24 26.94 24.64 21.47 48.84 46.97 44.73 43.24 42.47
- 65.16 64.48 57.04 56.44 51.56 46.49 44.03 39.70 59.59 58.49 57.35 56.30 54.71
- 65.78 65.17 57.49 56.39 50.35 45.92 43.73 37.66 56.48 55.12 53.98 52.79 53.40
- 64.98 64.22 58.03 56.84 45.23 40.77 38.29 33.88 54.49 52.64 51.21 50.04 50.88
65.99 65.29 57.50 56.90 53.44 48.89 46.41 42.37 60.13 58.93 57.40 56.26 55.79
A.1 Performance comparison between SS3D and SS3D++
We conduct the experiments on the Waymo dataset to compare SS3D++ with SS3D [23]. Tab. XII shows the comparison results, from which it can be seen that the performance of SS3D++ on the Waymo dataset has significantly improved compared to SS3D. SS3D++ shows the highest average performance across all categories and levels, suggesting it is the most robust model among those tested.
TABLE XII: Performance comparison with SS3D on the Waymo Open Dataset with 202 validation sequences for the 3D detection. †: with centerhead.
Methods Data Veh. (LEVEL 1) Veh. (LEVEL 2) Ped. (LEVEL 1) Ped. (LEVEL 2) Cyc. (LEVEL 1) Cyc. (LEVEL 2) Avg.
mAP mAPH mAP mAPH mAP mAPH mAP mAPH mAP mAPH mAP mAPH
CenterPoint [78] Sparse (2.2%) 32.98 32.33 28.82 28.24 20.22 16.91 17.30 14.47 28.84 27.90 27.74 26.84 25.22
SS3D [23] Sparse (2.2%) 49.33 48.40 41.93 41.03 35.74 30.98 29.57 25.88 40.36 39.19 38.24 37.54 38.18
SS3D++ Sparse (2.2%) 63.49 62.93 55.30 54.81 48.93 43.89 41.53 37.22 51.87 50.81 50.07 49.04 50.82
†Voxel-RCNN [5] Sparse (2.2%) 47.83 46.32 41.83 40.51 25.88 20.73 21.78 17.44 50.29 48.48 48.38 46.65 38.01
SS3D [23] Sparse (2.2%) 50.74 49.38 45.11 44.47 40.56 36.01 37.94 34.03 53.27 51.35 50.88 48.76 45.20
SS3D++ Sparse (2.2%) 65.99 65.29 57.50 56.90 53.44 48.89 46.41 42.37 60.13 58.93 57.40 56.26 55.79
Appendix B Computational Overhead Comparison
To analysis the trade off between SS3D and SS3D++, we compare the computational overhead of SS3D and SS3D++ on both the Waymo and KITTI datasets. All experiments are conducted on the Tesla A100 (40GB) GPU, paired with an Intel(R) Xeon(R) Gold 5318Y @ 2.10GHz CPU. The software environment includes PyTorch version 1.10.0 and CUDA version 11.3. As shown in Tab. XIII, the computational overhead comparison reveals that SS3D++ incurs a slight increase in training time and computational cost compared to SS3D. Notably, for the Waymo dataset, SS3D++ demonstrates a 23.4% improvement in average performance, increasing from 45.20 to 55.79. However, this comes with a 16.6% increase in training time and a 9.1% increase in memory usage. Therefore, the substantial performance gains achieved by SS3D++ make the additional overhead well worth the improvement in results, indicating that this is an essential improvement.
TABLE XIII: Comparison of computational overhead between SS3D and SS3D++.
Metric Waymo KITTI
SS3D SS3D++ Change (%) SS3D SS3D++ Change (%)
Training Time (hours) 8.4h 9.8h $\Delta$16.6% 15.9 16.8h $\Delta$5.6%
Memory Usage (GB) 60.3GB 65.8GB $\Delta$9.1% 30.7GB 32.3GB $\Delta$5.2%
Training Resources 4 $\times$ A100 4 $\times$ A100 $\Delta$0% 2 $\times$ A100 2 $\times$ A100 $\Delta$0%
Backbone Detector Voxel-RCNN Voxel-RCNN $\Delta$0% PV-RCNN PV-RCNN $\Delta$0%
Performance Avg. 45.20 55.79 $\Delta$23.4% 80.87 82.89 $\Delta$2.5%
Appendix C Annotation Analysis for the proposed sparsely-supervised method
Figure 16: (a) Illustration of the annotation process for objects with varying levels of difficulty: on the left, priority is given to annotating “Easy” objects, while on the right, priority is given to annotating “Hard” objects. Besides, with the “Random” setting, we can validate the impact of which box to annotate. (b) The process of introducing noise annotations by perturbing accurate annotations.
In our sparsely-annotated framework, we just annotate one 3D object per scene. A natural question arises: which object to annotate, and how does the accuracy of the annotations impact the results?
On the one hand, most annotators may be more inclined to choose the easiest object to annotate in practice, e.g., no occluded objects or high-density objects. To validate the influence of which box to annotate, we design the following three different annotated strategies: (1) randomly choose one object from the scene to annotate as default; (2) prioritize choosing one ‘easy’ (close and no truncated) object from the scene; (3) prioritize choosing one ‘hard’ (remote and truncated) object from the scene. Different annotated strategies are shown in Fig. 16 (a). Besides, we repeat the experiment 3 times and average the results to eliminate the influence of experimental randomness. From Tab. XIV, we can observe that the hardness of annotated objects has a significant impact on the performance of the fully supervised detector, in which annotated strategy (3) severely deteriorates the detectors. However, by employing the proposed SS3D++ method, regardless of different annotated strategies, such annotation bias can be almost eliminated. Especially, the SS3D++ can achieve on-par performance compared to detectors trained with fully-supervised data for all cases. This shows that our method could perform robustly with different annotated strategies, which highlights the capability of our method to effectively address such low-quality annotations and further substantiates its potential practical value.
TABLE XIV: Annotation analysis of three different sparsely-annotated forms. “(1) (2) (3)” denotes different sparse data settings: (1) randomly choose one object from the scene to annotate as default; (2) prioritize choosing one “easy” object from the scene; (3) prioritize choosing one “hard” object from the scene.
Method Data Car - 3D Detection Car - BEV Detection Avg.
Easy Mod Hard Easy Mod Hard
Voxel-RCNN Full 92.38 85.29 82.86 95.52 91.25 88.99 89.38
Voxel-RCNN (1) Sparse 72.97 63.05 60.23 81.64 74.91 71.93 70.78
Ours (SS3D++) (1) Sparse 93.50 84.36 82.64 96.63 91.31 89.02 89.57
Improvements - +20.53 +21.31 +22.41 +14.99 +16.40 +17.09 +18.78
Voxel-RCNN (2) Sparse 88.86 63.04 47.90 93.61 70.82 52.77 69.50
Ours (SS3D++) (2) Sparse 92.89 83.46 79.81 96.40 91.63 86.73 88.48
Improvements - +4.03 +20.42 +31.91 +2.79 +20.81 +33.96 +18.98
Voxel-RCNN (3) Sparse 29.15 42.92 49.62 35.81 53.83 60.90 45.37
Ours (SS3D++) (3) Sparse 94.24 84.11 82.44 96.65 91.38 88.99 89.63
Improvements - +65.09 +41.19 +32.82 +60.84 +37.55 +28.09 +44.26
On the other hand, inaccurate annotations introduce noise to the annotations, highlighting the inherent challenges of the machine learning task. This label noise can significantly impede model training, limiting the model’s ability to learn effectively. To assess the impact of annotation accuracy, as shown in Fig. 16 (b), we randomly perturb some of the annotated boxes, ensuring that the IoU of the perturbed boxes with the original ground truth boxes fall within the range of $[0.45,0.55]$. Additionally, we set two perturbation ratios: 20%, and 50%, indicating the proportion of boxes that are perturbed. The experimental results, as presented in Tab. XV, indicate that varying ratios of perturbation adversely affect both the fully supervised method and our approach. As the perturbation ratio increases, the detrimental effects become increasingly pronounced. However, in comparison to the fully supervised method, our approach exhibits greater robustness, consistently outperforming it across different ratios of perturbation.
TABLE XV: Annotation analysis of different perturbation ratios.
Method Data Perturbation Car - 3D Detection Car - BEV Detection Avg.
Ratio Easy Mod Hard Easy Mod Hard
Voxel-RCNN Full 0% 92.38 85.29 82.86 95.52 91.25 88.99 89.38
Voxel-RCNN Sparse 0% 75.55 64.67 62.43 83.34 76.75 73.50 72.71
Ours (SS3D++) Sparse 0% 93.59 83.83 82.79 96.71 91.70 89.26 89.65
Improvements - - +1.21 -1.46 -0.07 +1.19 +0.45 +0.27 +0.27
Voxel-RCNN Full 20% 92.07 82.80 80.11 95.46 90.79 87.60 88.14
Voxel-RCNN Sparse 20% 74.98 65.65 61.83 82.98 75.52 72.25 72.20
Ours (SS3D++) Sparse 20% 92.82 82.86 79.99 96.33 91.00 88.44 88.57
Improvements - - +0.75 +0.06 -0.12 +0.87 +0.21 +0.84 +0.44
Voxel-RCNN Full 50% 92.17 82.39 79.83 95.64 88.56 86.98 87.60
Voxel-RCNN Sparse 50% 73.17 61.79 58.33 81.17 74.53 72.06 70.18
Ours (SS3D++) Sparse 50% 92.85 81.89 79.03 96.12 89.99 87.71 87.93
Improvements - - +0.68 -0.50 -0.80 +0.48 +1.43 +0.73 +0.34
It can be observed that training a detector with the original fully-supervised network under different data annotation conditions yields varying results. For instance, training the network with prioritize keeping easy objects results in higher performance for the ”Easy” difficulty level but lower performance for the ”Hard” difficulty level. Conversely, training the network with prioritize keeping difficult objects leads to lower performance for the ”Easy” difficulty level but higher performance for the ”Hard” difficulty level. However, after embedding the network into the proposed framework, regardless of the annotation condition, SS3D++ still achieves on-par performance comparing to fully-supervised method under three difficulty, indicating practical value.
To sum up, the two experiments mentioned above validate the robustness of our method in terms of annotation accuracy and annotation randomness, further highlighting the immense potential of our method in practical applications.
Appendix D Additional Analysis of Visualization Comparison
D.1 Comparison on the KITTI Dataset
The detection results of PointRCNN [44] and PartA2 [45] are shown above the dashed line in Fig. 18. As can be seen, the proposed SS3D++ method can obviously improve the detection performance and overcome the various error cases with different detector, which further demonstrates the superiority of the detector-agnostic property.
D.2 Comparison on the Waymo Open Dataset
Appendix E Additional Analysis of Pseudo Instances
Appendix F Additional Analysis of Learning from Unlabeled Training Scenes in An Online Manner
Algorithm 4 Algorithm of Learning from Streaming Unlabeled Scenes.